in a certain factory producing cycle tyres there is a small chance of 1 in 500 tyres to be defective. the tyres are supplied in lots of 10 . using Poisson distribution, calculate the approximate number of lots containing no defective one defective and two defective tyres respectively in consignment of 10,000 lots

1
Expert's answer
2021-11-01T20:01:36-0400

Binomial distribution with parameters "n = 10" and "p =1\/500= 0.002" can be approximated using Poisson with parameter "\lambda = np=10(0.002) = 0.02."

(a)


"P(X=0)=\dfrac{e^{-0.02}\cdot(0.02)^0}{0!}=e^{-0.02}"

"\approx0.98019867"

"0.98019867(10000)=9802"

(b)


"P(X=1)=\dfrac{e^{-0.02}\cdot(0.02)^1}{1!}=0.02e^{-0.02}"

"\approx0.01960397"

"0.01960397(10000)=196"

(c)


"P(X=2)=\dfrac{e^{-0.02}\cdot(0.02)^2}{2!}=0.0002e^{-0.02}"

"\approx0.00019604"

"0.00019604(10000)=2"